Fermi National Laboratory


Accelerator Update for April 15, 2005 - April 18, 2005

Friday April 15
The day shift began with Operations monitoring store 4098, the Recycler stash, and the Antiproton Source (Pbar) stack, and with experts investigating a magnet problem in the P1 or P2 transfer line.

A Pbar expert began a study period at 8:50 AM.

A Main Injector (MI) expert began a study period at 10:45 AM.

Operations terminated store 4098 around 11:33 AM. Operations then began preparing the Tevatron (TeV) for an access into F sector to inspect the P1 and P2 transfer lines, which are in F0 and F1 respectively.

Accelerator Overview

The TeV tunnel accesses began at 12:28 PM.

Experts in the P2 transfer line reported at 2:23 PM finding a quadrupole magnet with a turn-to-turn short. They began the process of replacing the magnet.

Pbar experts accessed AP0 to conduct maintenance at 4 PM.

Operators prepared the Linac for an access at 7:14 PM.

Saturday April 16
The midnight shift began with Operations monitoring the stack and stash, with Booster experts conducting studies, and with the P2 magnet change continuing.

An expert reported at 8:34 AM that the P2 quadrupole magnet (QF15) was installed and was ready for survey. The cryo system cool down for F sector began.

The Cryo Coordinator reported at 1:16 PM that the F sector cryo system was ready to go.

Operations established beam to MiniBooNE around 4 PM.

Operation put the TeV into recovery mode at 5:56 PM.

Operations established beam to the MI dump at 6:21 PM and beam to the Accumulator five minutes later.

Operations began ramping the TeV at 6:42 PM.

Operations put the TeV through a dry squeeze at 6:58 PM

Operations established beam to the Meson MCenter experiment at 10:11 PM.

Sunday April 17
The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 12:44 AM, a Booster RF station (BRF16) tripped off. It reset okay.

At 1:38 AM, operators reported a problem with the Linac beam at Tank #1. This required them to lower the beam intensity to MiniBooNE and to tune Linac.

Operations put the TeV into shot setup at 1:49 AM.

At 2:30 AM, BRF16 tripped off. It reset okay.

At 3:04 AM, the P1 transfer line suffered a BPM problem that held up tuning the line.

Operations put the TeV through a wet squeeze around 4 AM.

At 5:12 AM, BRF16 tripped off. It reset okay.

Operations established combination store 4100 at 6:19 AM with an initial luminosity of 98.14E30. Pbar resumed stacking.

MCenter resumed taking beam at 6:28 AM.

At 6:39 AM, BRF4 tripped off for the first of two times in an hour. It reset okay both times.

At 8:41 AM, BRF16 tripped off for the first of three times in a little over two hours. It reset okay all three times.

At 9:37 AM, an Accumulator quadrupole (A:LQ) power supply tripped off, which caused the loss of the stack. Operators called in an expert.

The Accumulator expert reported at 12:16 PM that there was a regulator problem and that he needed to call in more help.

At 12:42 PM, two Booster RF stations (BRF7 & 8) tripped off due to a vacuum burst. Everything reset okay.

Accumulator experts reported at 2:10 PM that they could not find an A:LQ regulation problem. The swapped the circuitry with a different device.

Pbar resumed stacking at 5:51 PM.

At 5:53 PM, MiniBooNE experts asked to have their beam halted while they resolved a data acquisition problem.

At 5:53 PM, BRF16 tripped off for the first of many times in a couple of hours. It reset okay each time.

Operations reestablished beam to MiniBooNE at 6 PM.

At 8:47 PM, BRF15 tripped off. It reset okay.

At 9:43 PM, BRF16 tripped off multiple times. Operators bypassed the station.

At 10:37 PM, MIRF14 tripped off for the first of three times in ninety minutes. It reset okay each time.

Monday April 18
The midnight shift began with Operations monitoring store 4100, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 1:54 AM, BRF15 tripped off. It reset okay.

At 4:22 AM, MIRF14 tripped off for the first of two times in eight minutes. It reset okay each time.

Operations managed to get BRF16 to reset at 4:29 AM.

Operations put the Recycler into shot setup to stash antiprotons at 6:03 AM.

At 6:25 AM, BRF16 tripped off. It reset okay.

Recycler stashing ended around 7:16 AM. Pbar resumed stacking and MCenter resumed taking beam.

The Plan for Monday
The plan for today is to terminate store 4100 around 11 AM. The Run Coordinator will allow TeV experts to install some new octupole curves and then test them with a study period. Operations will establish a new store later this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/18/2005   email Fermilab